OpenStudy (anonymous):

trace the curve/ graph r(3-sintheta)=a

OpenStudy (hunus):

\[\Large r = a*\frac{1}{3-\sin(\theta)}\] Do you have any idea where to start?

OpenStudy (anonymous):

Let 'a' be your divisions on one axis and then steadily increase theta and compute the values plotting them (as precisely as possible) for one cycle of 2 pi.